“The series expansion of the function cosx in powers of x contains all even 

 powers of x only.” True/ False ?



1
Expert's answer
2021-10-06T10:07:25-0400
"\cos x=1-\dfrac{x^2}{2!}+\dfrac{x^4}{4!}-\dfrac{x^6}{6!}+...=\displaystyle\sum_{n=0}^{\infin}(-1)^n\dfrac{x^{2n}}{(2n)!},"

for all "x."


“The series expansion of the function cosx in powers of x contains all even 

 powers of x only.”  True.